OpenStudy (anonymous):

Solve each problem by using the quadratic formula. Write solutions in simplest radical form. 2x2 + 2x – 1 = 0

OpenStudy (anonymous):

@cupcake111

OpenStudy (anonymous):

you know this is on google... your better off looking up easy answers on google than wasting your time waiting on here for people to answer. i found lots of links about this on google.

OpenStudy (anonymous):

https://answers.yahoo.com/question/index?qid=20100506122419AA60RMm #6 on there

OpenStudy (anonymous):

nevermind, that one is slightly different , it has -2x, and yours has positive 2x.

OpenStudy (anonymous):

olve each problem by using the quadratic formula. Write solutions in simplest radical form. 2x^2 - 2x - 1 = 0 This is in the form ax^2 + bx + c a = 2, b = 2, c = -1 Quad formula is (-b +/- [sqrt(b^2-4ac)])/2a Plug in for that.

OpenStudy (anonymous):

\(\Large x=\frac{-2\pm \sqrt{2^2-4\times2\times(-1)}}{2\times2}=\frac{-1\pm \sqrt{3}}{2}\)

OpenStudy (anonymous):

correct!